1/1

Karl Friedrich Müchler,

Down deep within the cellar

tradukita de John Oxenford

Down deep within the cellar, here Against a butt I rest, Sir, My heart of ev’ry care I clear, And swear I’ll have the best, Sir.

The cellar man taps any cask, I choose by nod or winking; He fills the glass, he loves his task As I love drinking, drinking!

The demon thirst I could withstand, Nay, if I could, would kill it; I therefor take my glass in hand And with good liquor fill it.

The world then wears a hue so gay, It nearly sets me blinking; All evil thoughts I scare away, While drinking, drinking, drinking!

Bu oh, at ev‘ry glass I drain, That ragging thirst increases, To drown it, tiplers strive in vain, It never, never ceases.

Yet where’s the odds if down I fall? E’en now I feel a sinking; I do not harm, wish well to all, While drinking, drinking, drinking!

Traduko de la Germana poemo “Im tiefen Keller sitz ich hier” de Karl Friedrich Müchler (∗1763-09-02 – †1857-01-12) en la Anglan de John Oxenford.

Arg-944-1898 (2013-12-04 10:09:38)

Pri la tradukinto John Oxenford vidu la retejon http: // en. wikipedia. org/ wiki/ John_Oxenford .

http://www.poezio.net/version?poem-id=944&version-id=,1898